### v1

@@ Abstract
Quantum metrology offers superior measurement precision compared to classical counterparts. Traditionally, achieving this enhancement involves employing multi-particle entangled quantum states, necessitating state preparation and detection schemes that are practically challenging. In this article, we present a versatile and on-demand protocol for deterministic parameter estimation that leverages two state-transfer operations on a single bosonic mode. Specifically, we demonstrate this protocol in the context of phase and amplitude estimation using the superposition of coherent states in the bosonic circuit quantum electrodynamics (cQED) platform. With low photon numbers of up to 1.76, we achieve quantum-enhanced precision approaching the Heisenberg scaling, reaching a metrological gain of $7.5(6)$ dB and $9.3(5)$ dB respectively for phase and amplitude estimation. We show that the gain or sensitivity range can be further enhanced on the fly by tailoring the input states based on specific system constraints. Our protocol can be adapted to further optimize the desired figures of merit using tailored quantum states or operations, not only for bosonic cQED hardware but also readily extensible to other continuous-variable platforms. The realisation of this versatile and effective scheme affords a promising path towards practical quantum-enhanced sensing.
@@
One of the earliest and most canonical examples is quantum-enhanced phase estimation in a two-mode interferometer [14, 15]. The use of coherent states (CS) defines the standard quantum limit (SQL) with precision $\Delta\theta=1/\sqrt{N}$, where $\Delta\theta$ is the standard deviation of the phase and $N$ is the number of particles or excitations. Using entanglement as the resource state in this configuration, the precision scales as $\Delta\theta=1/N$ (known as the Heisenberg limit), which affords a $\sqrt{N}$ improvement over the SQL. This enhanced precision has been demonstrated in many experiments using entangled multipartite systems [16, 17, 18, 19, 20]. Practically, however, both the preparation of these non-local states and the implementation of correlated or conditional measurements can be challenging. Furthermore, preparing such quantum states often demands post-selection or heralding, which undermines the quantum enhancement [21, 22] over deterministically created coherent states.
@@
Figure 1: Universal quantum metrology with a single bosonic mode. The approach consists of three fundamental steps: preparation of input state $|\psi\rangle$; distortion from the to-be-estimated process, potentially induced by stray fields, which effectively encode the parameter of interest $\beta$ into the state $|\psi(\beta)\rangle$; and information retrieval by projection onto the initial state, depicted by the shaded region.
@@
In light of these challenges, there has been growing interest in quantum metrology using non-entangled states [23]. For instance, Fock states are used to perform phase-insensitive measurements with enhanced precision due to their nonclassical features [24, 25]. Quantum-enhanced phase estimation has also been shown using displaced Fock states or their superposition [26, 27, 28]. This handily bypasses the demanding non-local operations, which generally suppress the quantum advantage under even slight decoherence. However, the existing single-mode schemes are system-specific, as they rely on highly-specialised input states or quantum gates.
@@
Figure 2: Parameter estimation procedure with bosonic cQED. (a) Implementation of the protocol in cQED. A cavity stores a bosonic mode, coupled to a transmon qubit. The readout resonator provides single-shot measurements of the qubit state. The protocol includes a state-transfer operation $\hat{O}_{1}$, a process $\hat{U}(\beta)$, and a projection operation involving a state-transfer operation $\hat{O}_{2}$, qubit measurement $p_{g}$, and cavity measurement $p_{0|g}$. (b) Illustration of phase estimation using SCS. The shaded area visualises the overlap between the distorted and initial state $|\langle\psi|\psi(\theta)\rangle|^{2}$. (c) The FI for SCS (solid blue) and CS (solid purple), calculated using Eq. (1), where $p$ is from the polynomial fit (dotted curves) of the experimental $p_{g0}(\theta)$ (markers) in the inset. The shaded area with the corresponding color indicates the standard deviation error from bootstrapping. Bandwidth is defined as the range on which the FI of SCS is above the maximum FI ($\rm{FI}_{max}$) of CS. The grey area represents the region where we have a quantum advantage compared to CS and $\tilde{\rm{FI}}$ is its corresponding average FI. Dashed curves show simulation. (d) The $\rm{FI}_{max}$ for SCS (blue) and CS (purple) are plotted as markers for states with different photon numbers $\overline{n}$. The errors are standard deviation from bootstrapping. Dashed curves are simulation and dotted curves are fitting: $3.92(9)\overline{n}^{1.71(1)}+0.15(4)$ for SCS and $2.62(6)\overline{n}^{0.99(1)}+0.00(4)$ for CS. The solid pink curve is the ideal QFI of SCS.
@@
Here, we demonstrate a versatile and on-demand protocol for deterministic parameter estimation using a single bosonic mode. With standard bosonic cQED hardware, we show quantum-enhanced performance for both phase and amplitude estimation, achieving state-of-the-art metrological gain in the low photon number regime. We present a comprehensive analysis of key metrological figures of merit, including the sensitivity bandwidth, average sensitivity within the bandwidth, and adjustment of the position of maximum sensitivity. Importantly, we show that our implementation allows these key figures to be optimally traded on the fly to best suit the intended metrological goal by simply varying a few programmable pulse parameters in the protocol. Our strategy is universal across different physical systems and can be applied to other physical implementations of bosonic modes, such as optical resonators [29], trapped ions [30], neutral atoms [31], and mechanical resonators [32].
@@
Figure 3: Comprehensive analysis for phase estimation and quantum-enhanced precision. (a) The average FI ($\tilde{\rm{FI}}$) of SCS within the bandwidth and $\theta$ offset at which FI is maximum vs $\overline{n}$. The dotted curve is a fit: $2.7(3)\overline{n}^{2.0(1)}+0.3(1)$. (b) The same protocol as in Fig. 2(c) with initial states prepared with offset in $\theta$. As the protocol is symmetrical with respect to $\theta$, maximum FI at $\theta=0$ can be obtained by a positive/negative offset of about $0.2-0.25$ rad. (c) The phase precision vs $\overline{n}$ for both SCS (blue circles) and CS (purple diamonds). The solid lines are different theoretical Heisenberg limits and SQL. The dotted lines are linear fit: $\log(\Delta\theta)=-0.915(2)\log(\overline{n})-0.987(1)$ for SCS and $\log(\Delta\theta)=-0.464(1)\log(\overline{n})-0.508(1)$ for CS. The highest metrological gain for SCS compared to CS with the same photon number is $7.5(6)$ dB.
@@
Conceptually, our methodology is akin to shining lights on an object under the influence of some external perturbations and comparing the resulting shadow with the original unperturbed object, as illustrated in Fig. 1(a). The shadow represents the projection of the distorted state $|\psi(\beta)\rangle=\hat{U}(\beta)|\psi\rangle$ onto the initial state $|\psi\rangle$, written as $p(\beta)=|\langle\psi|\psi(\beta)\rangle|^{2}$ where $\beta$ is the to-be-estimated parameter, $\hat{U}(\beta)=\exp(i\hat{B}\beta)$, and we shall refer to $\hat{B}$ as the process generator. The measurement outcome $p(\beta)$ provides an optimal precision for estimating $\beta$ [33], as indicated by the Fisher information (FI)
@@
which provides a lower asymptotic bound on the achievable precision, i.e., $\Delta\beta\geq 1/\sqrt{F}$, known as the Cramér-Rao bound [34]. The optimality transpires as the FI saturates its maximum value given by the quantum Fisher information (QFI) $F_{q}=4(\Delta\hat{B})^{2}$, where $(\Delta\hat{B})^{2}=\langle\psi|\hat{B}^{2}|\psi\rangle-\langle\psi|\hat{B}|\psi\rangle^{2}$. In addition to having an optimal measurement scheme, it is also important to choose the input states strategically for optimal metrological performance, i.e., those with high variance $(\Delta\hat{B})^{2}$.
@@
We implement this generalised metrology scheme on standard cQED hardware as depicted in Fig. 2(a). Initially, a desired quantum state is prepared with a state-transfer unitary $\hat{O}_{1}$, i.e., $|g\psi\rangle=\hat{O}_{1}|g0\rangle$. After the system undergoes a process $\hat{U}(\beta)$, a projection measurement is performed with a unitary $\hat{O}_{2}$ and qubit-cavity measurements. The operation $\hat{O}_{2}$ can be done with any state-transfer unitary such that $|g0\rangle=\hat{O}_{2}|g\psi\rangle$. Finally, standard measurements are performed to obtain the probabilities of the qubit in the ground state $p_{g}$ and cavity in vacuum $p_{0|g}$, which is achieved by driving the qubit with a vacuum-selective $\pi$-pulse and measuring the qubit in the excited state. The projection yields the probability $p_{g0}=|\langle\psi|\hat{U}(\beta)|\psi\rangle|^{2}$ [33]. An important advantage of our protocol is that any two operations $\hat{O}_{1}$ and $\hat{O}_{2}$ realising the required state transfers are sufficient. This minimalistic approach makes it highly adaptable to a variety of hardware parameters and readily extensible to different physical systems with their preferred native gates. For instance, in cQED platforms featuring small qubit-oscillator dispersive coupling, $\hat{O}_{1,2}$ can be realised using echoed conditional displacement (ECD) gates [35, 36]; whereas devices with more significant dispersive coupling can leverage on gradient ascent pulse engineering (GRAPE) [37, 38, 39] or selective number-dependent arbitrary phase gates [40]. Beyond cQED, the state transfer operations can be achieved through coherent coupling of the internal-motional states of trapped ions/atoms [41, 27], ECD gates in an atom-light system in cavity-QED [42], or swap gates in qubit-mechanical resonator system in quantum acoustics [43].
@@
We first demonstrate this universal protocol in the context of phase estimation in cQED. The target process is enacted via the dispersive qubit-cavity coupling with Hamiltonian $H_{\text{int}}/\hbar=(\chi/2)\hat{\sigma}_{z}\hat{a}^{\dagger}\hat{a}$, where $\chi$ is the coupling strength, $\hat{\sigma}_{z}$ the $z$-Pauli matrix, and $\hat{a}$($\hat{a}^{\dagger}$) is the annihilation (creation) operator of the bosonic cavity. For a qubit in the ground state $|g\rangle$, the dynamics realises a unitary for the cavity $\hat{U}(\theta)=\exp(i\hat{a}^{\dagger}\hat{a}\theta)$, with $\theta=\chi t/2$. As a result, the process generator is given by $\hat{n}=\hat{a}^{\dagger}\hat{a}$, which means useful states here are those with high variance in photon number $(\Delta\hat{n})^{2}$.
@@
Theoretically, we investigate states with high photon number variances such as (listed from lowest to highest) superposition of Fock states, superposition of coherent states (SCS), and squeezed vacuum. Experimentally, we find SCS to be the most robust against device imperfections such as self-Kerr and offer the highest sensitivity in our system [33]. Therefore, we shall primarily focus on SCS, given by $\mathcal{N}_{\alpha}(|0\rangle+|\alpha\rangle)$ (in short SCS($\alpha$)), with the normalisation factor $\mathcal{N}_{\alpha}$, to demonstrate the efficacy of our strategy. The intuition behind the performance of SCS for phase estimation is understood by visualizing the projection $p_{g0}=|\langle\psi|\hat{U}(\theta)|\psi\rangle|^{2}$ in phase space, see Fig. 2(b). With this illustration in mind, we note two important observations. Firstly, as $\theta$ increases, the probability $p_{g0}(\theta)$ changes from $1$ to a constant value, where the fringes do not overlap and the probability is insensitive to a further increase in $\theta$. Secondly, the use of higher $\alpha$ (further right position of the right blob in Fig. 2(b)) will reduce the overlap more sharply at lower $\theta$, making the detection more sensitive to $\theta$ at the cost of a smaller range of sensitivity.
@@
Experimentally, we chose to implement the state transfer operations $\hat{O}_{1,2}$ using GRAPE, which is a convenient and widely used tool in cQED [37, 38, 39, 28]. However, this choice is not unique, e.g., in the weak dispersive regime, one may opt to use ECD gates to create SCS [44]. By executing the gate sequence in Fig. 2(a), we directly obtain the probability $p_{g0}(\theta)$. An exemplary plot comparing the use of SCS($\alpha=2$) and CS with the same average photon number $\overline{n}=\langle\hat{a}^{\dagger}\hat{a}\rangle$, is presented in Fig. 2(c) inset. Each single-shot measurement is one bit of information and we perform $1000$ repetitions for each point to obtain the statistics. The experimental data shows good agreement with simulations (corresponding dashed curves), which take into account decoherences and measurement imperfections [33]. It is apparent from the probability that the SCS is more sensitive to the change in $\theta$ than the CS.
@@
To quantitatively examine the performance of our implementation, we compute the FI in Eq. (1), where the corresponding $p$ is a smooth function obtained from the polynomial fit of $p_{g0}(\theta)$. The resulting FI for SCS and CS are plotted in Fig. 2(c) as solid blue and purple curves, respectively. As expected, the maximum FI of SCS beats that of the CS at the cost of the sensitivity range. Furthermore, figure 2(d) shows the maximum FI for SCS (blue circles) and CS (purple diamonds) for states with different photon numbers. We note that decoherence and experimental imperfections hinder our results from reaching the ideal QFI of SCS (pink solid curve), as by simulations (corresponding dashed curves) with real device parameters. However, despite that, second-order polynomial fits (corresponding dotted curves) show that the $\text{FI}_{\text{max}}$ of SCS is nearly quadratic, whereas that of the CS is linear, which indicates a clear $\sqrt{\overline{n}}$ quantum enhancement in the estimation precision.

### v2

@@ Abstract
Quantum metrology offers the potential to surpass its classical counterpart, pushing the boundaries of measurement precision toward the ultimate Heisenberg limit. This enhanced precision is normally attained by utilizing large squeezed states or multi-particle entangled quantum states, both of which are often challenging to implement and prone to decoherence in real quantum devices. In this work, we present a versatile and on-demand protocol for deterministic parameter estimation that leverages two efficient state-transfer operations on a single bosonic mode. Specifically, we demonstrate this protocol in the context of phase estimation using the superposition of coherent states in the bosonic circuit quantum electrodynamics (cQED) platform. With low average photon numbers of only up to 1.76, we achieve quantum-enhanced precision approaching the Heisenberg scaling, reaching a metrological gain of $7.5(6)$ dB. Importantly, we show that the gain or sensitivity range can be further enhanced on the fly by tailoring the input states, with different superposition weights, based on specific system constraints. The realization of this versatile and efficient scheme affords a promising path towards practical quantum-enhanced sensing, not only for bosonic cQED hardware but also readily extensible to other continuous-variable platforms.
I INTRODUCTION
@@
One of the earliest and most canonical examples is quantum-enhanced phase estimation in a two-mode interferometer [14, 15]. The use of coherent states (CS) defines the standard quantum limit (SQL) with precision $\Delta\theta\propto 1/\sqrt{N}$, where $\Delta\theta$ is the standard deviation of the phase and $N$ is the number of particles or excitations. Using entanglement as the resource state in this configuration, the precision can scale as $\Delta\theta\propto 1/N$ (known as the Heisenberg limit), which affords a $\sqrt{N}$ improvement over the SQL. This enhanced precision has been demonstrated in many experiments using entangled multipartite systems [16, 17, 18, 19, 20]. Practically, however, both the preparation of these non-local states and the implementation of correlated or conditional measurements can be challenging. Furthermore, preparing such quantum states often demands post-selection or heralding, which undermines the quantum enhancement [21, 22] over deterministically created coherent states.
@@
Figure 1: Generalized protocol for quantum metrology with a single bosonic mode. Our approach consists of three fundamental steps: preparation of input state $|\psi\rangle$; distortion from the to-be-estimated process, e.g. stray electromagnetic fields, which effectively encode the parameter of interest $\beta$ into the state $|\psi(\beta)\rangle$; and final information retrieval by a projection operation that extracts the overlap with the initial state, shown as the shaded region.
@@
In light of these challenges, there has been growing interest in quantum metrology using non-entangled states [23]. For instance, Fock states are used to perform phase-insensitive measurements with enhanced precision due to their nonclassical features [24, 25]. Quantum-enhanced phase estimation has also been shown using displaced Fock states or their superposition. This handily bypasses the demanding non-local operations, which generally suppress the quantum advantage under even slight decoherence. However, existing single-mode schemes generally necessitate highly specialized state-dependent quantum gates, which are often expensive to implement [26, 27, 28] or require non-deterministic operations [26].
@@
Figure 2: Parameter estimation procedure with bosonic cQED. (a) Implementation of the protocol in cQED. A cavity stores a bosonic mode, coupled to a transmon qubit that facilitates operations and measurements. The readout resonator provides single-shot measurements of the qubit state. The protocol includes a state-transfer operation $\hat{O}_{1}$, a process $\hat{U}(\beta)$, and a projection operation involving a state-transfer operation $\hat{O}_{2}$, qubit measurement $p_{g}$, and cavity measurement $p_{0|g}$. (b) Illustration of phase estimation using the SCS. The shaded area visualizes the overlap between the distorted and initial state $|\langle\psi|\psi(\theta)\rangle|^{2}$. (c) The FI for SCS (solid blue) and CS (solid purple), calculated using Eq. (1), where $p$ is from the polynomial fit (dotted curves) of the experimental $p_{g0}(\theta)$ (markers) in the inset. The shaded area with the corresponding color indicates the standard deviation error from bootstrapping. Dynamical range is defined as the range on which the FI of SCS is above the maximum FI ($\rm{FI}_{max}$) of CS. The grey area represents the region where we have a quantum advantage compared to CS and $\tilde{\rm{FI}}$ is its corresponding average FI. Dashed curves show simulation. (d) The $\rm{FI}_{max}$ for SCS (blue) and CS (purple) are plotted as markers for states with different average photon numbers $\overline{n}$. The errors are standard deviation from bootstrapping. Dashed curves are simulation and dotted curves are fitting in the range $\bar{n}=[0.07,1.76]$: $3.92(9)\overline{n}^{1.71(1)}+0.15(4)$ for SCS and $2.62(6)\overline{n}^{0.99(1)}+0.00(4)$ for CS. The solid pink curve is the ideal QFI of SCS.
@@
Here, we demonstrate a versatile and on-demand protocol for deterministic parameter estimation using a single mode with a standard bosonic cQED hardware. We show quantum-enhanced performance for phase as well as amplitude estimation using the superposition of coherent states (SCS), achieving state-of-the-art metrological gain in the low photon number regime, which minimizes the performance penalties due to hardware imperfections. We present a comprehensive analysis of key metrological figures of merit, including the dynamical range and demonstrate that our scheme can be adapted on the fly to optimize specific performance metrics. This is accomplished by varying a few programmable pulse parameters in the protocol to modify the superposition weights of the SCS. The efficiency and versatility of our implementation allow on-demand optimization of key performance metrics tailored to the metrological goal and specific hardware constraints. Finally, our strategy is adaptable across different physical systems and can be applied to other physical implementations of bosonic modes, such as optical resonators [29], trapped ions [30], neutral atoms [31], and mechanical resonators [32].
@@
Figure 3: Comprehensive analysis for phase estimation and quantum-enhanced precision. (a) The average FI ($\tilde{\rm{FI}}$) of SCS within the dynamical range and $\theta$ offset at which FI is maximum vs $\overline{n}$. The dotted curve is a fit: $2.7(3)\overline{n}^{2.0(1)}+0.3(1)$. (b) The same protocol as in Fig. 2(c) with initial states prepared with offset in $\theta$. As the protocol is symmetrical with respect to $\theta$, maximum FI at $\theta=0$ can be obtained by a positive/negative offset of about $0.2-0.25$ rad. (c) The phase precision vs $\overline{n}$ for both SCS (blue circles) and CS (purple diamonds). The solid lines represent theoretical Heisenberg scaling with different factors. The dotted lines are linear fit in the range $\bar{n}=[0.07,1.76]$: $\log(\Delta\theta)=-0.915(2)\log(\overline{n})-0.987(1)$ for SCS and $\log(\Delta\theta)=-0.464(1)\log(\overline{n})-0.508(1)$ for CS. The highest metrological gain for SCS compared to CS with the same photon number is $7.5(6)$ dB.
II THE PROTOCOL
@@
Conceptually, our methodology is akin to shining lights on an object under the influence of some external perturbations and comparing the resulting shadow with the original unperturbed object, as illustrated in Fig. 1. The shadow represents the projection of the distorted state $|\psi(\beta)\rangle=\hat{U}(\beta)|\psi\rangle$ onto the initial state $|\psi\rangle$, written as $p(\beta)=|\langle\psi|\psi(\beta)\rangle|^{2}$ where $\beta$ is the to-be-estimated parameter, $\hat{U}(\beta)=\exp(i\hat{B}\beta)$, and we shall refer to $\hat{B}$ as the process generator. The measurement outcome $p(\beta)$ provides an optimal precision for estimating $\beta$, see appendix A, as indicated by the Fisher information (FI)
@@
which provides a lower asymptotic bound on the achievable precision, i.e., $\Delta\beta\geq 1/\sqrt{F}$, known as the Cramér-Rao bound [33]. The optimality transpires as the FI saturates its maximum value given by the quantum Fisher information (QFI) $F_{q}=4(\Delta\hat{B})^{2}$, where $(\Delta\hat{B})^{2}=\langle\psi|\hat{B}^{2}|\psi\rangle-\langle\psi|\hat{B}|\psi\rangle^{2}$ denotes the variance.
In addition to having an effective measurement scheme, it is also important to choose the input states strategically for optimal metrological performance. An immediate approach would be to consider states with high variance such as the superposition of Fock states [27, 28], the superposition of coherent states (SCS), and squeezed vacuum, see Appendix C. In practice, the choice of state must also be considered with the realistic hardware constraints in mind. In the bosonic cQED platform, the key sources of imperfection include the decoherence of the ancillary qubit, which reduces the quality of the prepared states, as well as the cavity self-Kerr effect, which distorts the state in phase space. The effects of both are more severe for states with higher Fock state occupations. This motivates the key advantage of our choice of using the SCS, which has higher variance compared to the superposition of Fock states and at the same time maintains a lower photon occupation in the cavity compared to squeezed vacuum, see Appendix C. Thus, our strategy effectively maximizes the relevant metrological figures of merit in the presence of hardware imperfections.
We implement this generalized metrology scheme on a standard cQED hardware as depicted in Fig. 2(a). First, a desired quantum state is prepared with a state-transfer unitary $\hat{O}_{1}$, i.e., $|g\psi\rangle=\hat{O}_{1}|g0\rangle$, where $|g0\rangle=|g\rangle\otimes|0\rangle$ denotes the product state of the qubit-cavity system, with $|g\rangle$ the ground state of the qubit, and $|0\rangle$ the vacuum state of the cavity. After the system undergoes a process $\hat{U}(\beta)$, a projection measurement is performed with a unitary $\hat{O}_{2}$ and qubit-cavity measurements. The operation $\hat{O}_{2}$ can be done with any state-transfer unitary such that $|g0\rangle=\hat{O}_{2}|g\psi\rangle$. Finally, standard measurements are performed to obtain the probabilities of the qubit in the ground state $p_{g}$ and cavity in vacuum $p_{0|g}$, which is achieved by driving the qubit with a vacuum-selective $\pi$-pulse and measuring the qubit in the excited state. The projection yields the probability $p_{g0}=|\langle\psi|\hat{U}(\beta)|\psi\rangle|^{2}$, see Appendix A.
An important advantage of our protocol is that any two operations $\hat{O}_{1}$ and $\hat{O}_{2}$ capable of realizing the required state transfers are sufficient. The second operation $\hat{O}_{2}$ has been implemented with time reversal technique [34] ($\hat{O}_{2}=\hat{O}_{1}^{\dagger}$) or numerically optimized unitary [28], both of which have to guarantee the evolution of multiple states concurrently. In contrast, our approach only requires constructing a single state-transfer process corresponding to $|g0\rangle=\hat{O}_{2}|g\psi\rangle$. This provides an intuition on the appeal of our strategy, as there are infinitely many options for $\hat{O}_{2}$ that can perform the required state-transfer process whereas there is only one solution that satisfies $\hat{O}_{2}=\hat{O}_{1}^{\dagger}$. As a result, finding an appropriate experimental operation, either numerically or analytically, to enact the general $\hat{O}_{2}$, instead of $\hat{O}_{1}^{\dagger}$, is significantly more efficient.
@@
Such a minimalist approach makes our protocol highly adaptable to a variety of hardware parameters and readily extensible to different physical systems with their preferred native gates. For instance, in cQED platforms featuring small qubit-oscillator dispersive coupling, $\hat{O}_{1,2}$ can be realized using echoed conditional displacement (ECD) gates [35, 36]; whereas devices with more significant dispersive coupling can leverage on gradient ascent pulse engineering (GRAPE) [37, 38, 39] or selective number-dependent arbitrary phase gates [40]. Beyond cQED, the state transfer operations can be achieved through coherent coupling of the internal-motional states of trapped ions/atoms [41, 27], ECD gates in an atom-light system in cavity-QED [42], or SWAP gates in qubit-mechanical resonator system in quantum acoustics [43].
III PHASE ESTIMATION
@@
We demonstrate the performance of this scheme in the context of phase estimation. The target process is enacted via the dispersive qubit-cavity coupling with Hamiltonian $H_{\text{int}}/\hbar=(\chi/2)\hat{\sigma}_{z}\hat{a}^{\dagger}\hat{a}$, where $\chi$ is the coupling strength, $\hat{\sigma}_{z}$ the Pauli-$z$ matrix, and $\hat{a}$($\hat{a}^{\dagger}$) is the annihilation (creation) operator of the bosonic cavity. For a qubit in the ground state $|g\rangle$, the dynamics realizes a unitary for the cavity $\hat{U}(\theta)=\exp(i\hat{a}^{\dagger}\hat{a}\theta)$, with $\theta=\chi t/2$. As a result, the process generator is given by $\hat{n}=\hat{a}^{\dagger}\hat{a}$, which means useful states here are those with high variance in photon number $(\Delta\hat{n})^{2}$.
@@
Our preliminary experimental investigations confirm that the SCS indeed offers the most robust sensitivity (highest FI) compared to the superposition of Fock states and squeezed vacuum, see Appendix C. Therefore, we shall primarily focus on SCS, given by $\text{SCS}(\alpha)=\mathcal{N}_{\alpha}(|0\rangle+|\alpha\rangle)$, with the normalization factor $\mathcal{N}_{\alpha}$, to demonstrate the efficacy of our protocol. The SCS is simply a displaced cat state ($\mathcal{N}_{\alpha}(|-\alpha/2\rangle+|\alpha/2\rangle)$), and in the limit of large $\alpha$ the normalization factor is $\mathcal{N}_{\alpha}\approx 1/\sqrt{2}$. The intuition behind the performance of SCS for phase estimation is understood by visualizing the projection $p_{g0}=|\langle\psi|\hat{U}(\theta)|\psi\rangle|^{2}$ in phase space, see Fig. 2(b). With this illustration in mind, we note two important observations. Firstly, as $\theta$ increases, the probability $p_{g0}(\theta)$ changes from $1$ to a constant value, where the fringes do not overlap and the probability is insensitive to a further increase in $\theta$. Secondly, the use of higher $\alpha$ (further right position of the right blob in Fig. 2(b)) will reduce the overlap more sharply at lower $\theta$, making the detection more sensitive to $\theta$ at the cost of a smaller range of sensitivity.
@@
Experimentally, we chose to implement the state transfer operations $\hat{O}_{1,2}$ using GRAPE, which is a convenient and widely used tool in cQED [37, 38, 39, 28]. However, this choice is not unique, e.g., in the weak dispersive regime, one may opt to use ECD gates to create SCS [44]. By executing the gate sequence in Fig. 2(a), we directly obtain the probability $p_{g0}(\theta)$. An exemplary plot comparing the use of SCS($\alpha=2$) and CS with the same average photon number $\overline{n}=\langle\hat{a}^{\dagger}\hat{a}\rangle$, is presented in Fig. 2(c) inset. Each experimental point is an average of $1000$ single shot repetitions. The data shows good agreement with simulations (corresponding dashed curves), which take into account decoherences and measurement imperfections, see Appendix D.
@@
To quantitatively examine the performance of our implementation, we compute the FI in Eq. (1), where the corresponding $p$ is a smooth function obtained from the polynomial fit of $p_{g0}(\theta)$. The resulting FI for SCS and CS are plotted in Fig. 2(c) as solid blue and purple curves, respectively. As expected, the maximum FI of SCS beats that of the CS at the cost of the sensitivity range. Furthermore, figure 2(d) shows the maximum FI ($\text{FI}_{\text{max}}$) for SCS (blue circles) and CS (purple diamonds) for states with different average photon numbers. We note that decoherence and experimental imperfections hinder our results from reaching the ideal QFI of SCS (pink solid curve), as shown by simulations (corresponding dashed curves) with real device parameters. However, despite that, second-order polynomial fits in the range $\bar{n}=[0.07,1.76]$ (corresponding dotted curves) show that the $\text{FI}_{\text{max}}$ of SCS is nearly quadratic, whereas that of the CS is linear, which indicates a clear $\sqrt{\overline{n}}$ quantum enhancement in the estimation precision.
IV METROLOGICAL FIGURES OF MERIT